OpenStudy (anonymous):

how do you solve this equation -12+2v= -3(8v+4). Can anyone please show me the steps to solve these kinds of problems?

OpenStudy (anonymous):

-12+2v=-3(8v+4) -12+2v=-24v-12 +12 +12 2v=-24v+0 +24v 26v=0?

OpenStudy (anonymous):

i thought it would be that too, but then every problem would equal o and i don't think that's right. But tx

OpenStudy (anonymous):

theres a special term for when it equals 0 but i forget it